Sol Summary – March 9th

Crew 263 Sol Summary 09-03-2022

Sol: 16

Summary Title: Productive day

Author’s name: Cerise Cuny

Mission Status: Nominal

Sol Activity Summary: This morning, just as yesterday, we did not have a morning sports session altogether, as four of us had to preserve our energy for the EVA. The EVA went very well, we were able to do a very good amount of LIBS analysis. We were very impressed by the Summerville formation. Meanwhile, the three other crew members did some sport and prepared an excellent meal: rice salad with tuna and vegetables.

All day long we performed the ultrasound experiment for the French Space Agency. We are improving in terms of speed and quality of the acquisitions. We had diverse activities for other experiments such as Aquapad, Plant gravitropism, Spiruline growth with urine, etc. Lea and Marine had a lot of geological post-processing to do after this morning’s EVA and an exhaustive report.

Look Ahead Plan: Tomorrow will be a lighter day in terms of activities. In the morning, a very important EVA for Search & Rescue of Astronauts is planned in the North. Two crew members will go through their last session of a particularly long and exhausting human factors experiment.

Anomalies in work: None

Weather: Few clouds, except at noon (very cloudy). Light rain in the afternoon. Very high winds in the afternoon. Temperatures between -3 and 12°C.

Crew Physical Status: Very good.

EVA: EVA 20 in the morning. A suit’s battery died during EVA.

Reports to be filed: Sol Summary, EVA Report (n°20), EVA Request (n°21), Operations Report, Journalist Report, HSO Report, Astronomy Report.

Support Requested: None.

Research Report – March 9th

Name: Marine Prunier
Crew: 263
Date: 09 Mars 2022


Robotic Telescope Requested (choose one): MLC-RCOS16

Objects to be Imaged this Evening: None (New Mexico observatory closed for wind)

Images submitted with this report: /

Problems Encountered: The observations of LDN 1627 and M51 submitted yesterday on Skynet (8 Mars) were really blurry, a warning sign was displayed next to every observation file. The wind was less than a few Mph yesterday according to, I wonder what the issue was. When seeing that the first pictures were not very good, I canceled the rest of the observations (not to waste too many Skynet credits).


Solar Features Observed: /
Images submitted with this report: /
Problems Encountered: /

EVA Report – March 9th

Crew 263 EVA 20 Report 09-03-2022

EVA # 20

Author: Cerise Cuny

Purpose of EVA: Replacing the batteries for the atmospheric experiments, a second analysis of sulfate deposits using a handheld LIBS (Laser-Induced Breakdown Spectroscopy) analyser.

Start time: 08:09 am
End time: 12:19 pm

Narrative: We first checked the rovers’ batteries and usage duration, then drove with Opportunity (86.3h and 100%) and Curiosity (170.3h and 100%) to the atmospheric experiment site. Changing the batteries took only 5 minutes this time, everything went well. At 8:25 we left for Candor Chasma.

As planned, we stopped at the intersection of Galileo Road and Streambed connector at 8:40 (Opportunity 86% and Curiosity 88%). From there, the canyon was reached by foot. We found our path very easily thanks to the very precise instructions and pictures of yesterday’s EVA team.

Inside the canyon, we made several stops to proceed on pictures and LIBS analysis. We took a sampling of what we thought was a nodule for further analysis at the station. The drone was quite useful to evaluate which arms of the canyon were accessible by foot. We were able to reach outcrops that wouldn’t be reachable by a Martian rover.

We also tried to use it to spot interesting outcrops but at some point, the Summerville formation was everywhere around us, so it was easy to find relevant outcrops with our own eyes.

At 11:11, we u-turned and reached the rovers at 11:50, with several breaks to catch our breath. Elena’s suit went out of battery at 11:15 during our return, she struggled a bit with fog in the helmet. We entered the AirLock at 12:19 for depressurisation.

Opportunity: Usage Duration – 86.7 Battery Charge – 81%

Curiosity: Usage Duration – 170.7 Battery Charge – 81%

Destination: Zone G -final expected point of interest is located at coordinate 4251533.157 N 521043.495 E.

Participants: Cerise Cuny (EVA leader), Nicolas Wattelle, Mathéo Fouchet, Elena Lopez-Contreras Gonzalez

Road(s) and routes per MDRS Map: Entrance Road, Cow Dung Road, Galileo Road 1104.

Mode of travel: Rover for battery change and then until the entrance of Candor Chasma, By Foot in Candor Chasma.

Vehicles you will be using (If applicable): Curiosity, Opportunity

Operations Report – March 9th

Crew 263 Operations Report 09-03-2022

SOL: 16

Name of person filing report: Mathéo Fouchet

Non-nominal systems: Nothing to report

Notes on non-nominal systems: Nothing to report


Spirit rover used: No

Hours: 157.7

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Opportunity rover used: Yes

Hours: 86.7

Beginning charge: 100%

Ending charge: 81%

Currently charging: Yes

Curiosity rover used: Yes

Hours: 170.7

Beginning charge: 100%

Ending charge: 81%

Currently charging: Yes

Perseverance rover used: No

Hours: 228.6

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

General notes and comments: Nothing to report

Summary of Hab operations:

WATER USE: 17 gallons

Water (static tank): 245 gallons

Water (loft tank): 38 gallons

Water Meter: 01563570 units

Static to Loft Pump used: Yes

Static tank pipe heater: On

Static tank heater: On

Toilet tank emptied: Yes, and it is not even smelling when being emptied so that is good news

Summary of internet: Nothing to report

Summary of suits and radios: I glued the ring of the suit’s helmet as suggested by Scott yesterday. Then I put it in the RAM so that Shannon can take it and see what she can do about it. Also, during today’s EVA, suit #2 shut down after 3 hours of use. We were on our way back to the station so we could not do anything about it. I suppose it is quite normal for a long EVA due to the suit’s battery.

Summary of GreenHab operations: Nothing to report

Summary of ScienceDome operations:

Dual split: Off

SOC: 89% at 6:38pm

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Nothing to report

Journalist Report – March 9th

Sol 16 – Another day at the office

Author: Nicolas Wattelle

Another intense day on Mars, the kind of day we will remember once we will be back on Earth.

The program: the last session of ultrasound medical surveillance for everyone, finishing up the geology work in Candor Chasma for Mathéo, Elena, Cerise, and I, repairing the HF radio transceiver… and others.

Definitely, the ultrasound experiment might be our favourite. If you are a subject, you just lie on a comfortable bed, and for the next half hour, until you become what we call “first operator”. This latter is the actual person experimenting. This person has the probe in one hand, and the phone screen with the EchoFinder app on the other.

As I described for you in Sol 2’s report, it is then like a “game”, and to win you have to make a great picture. In the crew, we almost take that as a way to challenge each other, and as you might know, Elena is always the most valuable player… And once you have checked all organs asked, measured your performance, and given your feedback on the protocol given by CNES, you are done.

You become “operator 2”, to assist operator 1’s performance measurement. And because it is finally a time to do experiments together, to talk while changing probes, to share a moment, it is always a great time.

As I said before, the crew went back in Candor Chasma, to complete documentation of the zone, which began yesterday. But, because it was the same kind of “sporty EVA” as the previous one, this one was performed by the crew members that stayed inside yesterday. We had much fewer difficulties to go to the points of interest, for the simple reason that we were wisely briefed by the Martianauts who went there, not to make the same mistakes. And more, we flew the PARROT drone several times, which allowed us to have a clear vision of our environment.

We loved this outing. Being an operator in such conditions is unique, especially when it goes well: our “laser gun”, the LIBS Z-903, worked perfectly. The weather was good, the colours were magical. Take a look at the drone shot of the day!

During the last few days, we had trouble making our antenna setup work. Our last problem was on the feed line, where a transformer decided not to deliver enough voltage for a reason we ignore. We received advice from the Toulouse radio club to solve the problem: using a car battery charger and a car battery (that we usually use for atmospheric experiments). And it worked! You might see that it looks way more “handcrafted” than before, but we safely wired it: Safety first, even before science.

It is interesting seeing us working now, we feel like living in “normality” but in fact, we do not. The sequence of activities throughout the days, the station maintenance, the experiments, the EVAs … All of this makes our daily life, and we almost do not question it… It is just when taking a step back on it that you realise concretely. We might have become real Martians!

Supplemental Operations Report – March 9th

Supplemental Operations Report 9 Mar 2022

Name of person filing report: Shannon Rupert

Reason for Report: Routine (full week report, I missed last Sunday)

Non-nominal systems: Outpost water system, Sojourner

Action taken for non-nominal systems: I am emptying the water supply as much as possible, then I will shock it with bleach and refill the storage tank. I’m too big to get into the tank to clean it. I don’t know for sure that the water is bad but we have been having stomach issues and it was not cleaned at the beginning of the field season so this is just in case.

Sojourner’s batteries are not working properly and are losing charge much quicker than when I left in January. I will be checking and refilling all battery cells and see if that improves the power.

Generator: Still being manually run at night. Approximate hours 9 pm to 7 am on sunny days, slightly longer (with an earlier start time at night) if it is cloudy.

ScienceDome Dual Split: Off

Solar—Charge controllers off at night.

Solar— SOC is nominal.

Notes on power system: Nothing new to report. I will be working on the system this weekend, attempting to get the generator to charge the batteries at night.

Propane Reading, station tank – 64 %

Propane Reading, director tank— 70+ % (estimated at 80%)

Propane Reading, intern tank— 80 %

Propane Reading, generator— 70 %

Ethanol Free Gasoline – not known gallons.

Water (Outpost tank) – 75 gallons

Sojourner rover used: yes

Hours: 162.4

Beginning charge: 100

Ending charge: 100

Currently charging: Yes

Notes on rovers: Sojourner is losing charge far too quickly.

ATV’s Used: (Honda, 350.1, 350.2, 300): none

Reason for use: n/a

Oil Added? No.

ATV Fuel Used: 0 gallons

# Hours the ATVs were Used: 0 hours

Notes on ATVs:

HabCar used and why, where? Yes, to town for garbage runs and supplies

CrewCar used and why, where? Yes, to town

Luna used and why, where? Yes, to town

Campus-wide inspection, if action is taken, what and why? Nothing to report

General notes and comments: Nothing to report

Summary of general operations: Nothing to report

Summary of internet: Nothing new to report.

Summary of suits and radios: One helmet has damage and has been switched for the spare one at the Outpost.

Summary of Hab operations: Nothing to report

Summary of Outpost operations: Greywater tank valve needs replacement.

Summary of GreenHab operations: Nothing to report

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: Nothing to report

Summary of any observatory issues: We tried to open the observatory dome using an onsite internet connection but failed. We will attempt manually opening it next week.

Summary of health and safety issues: Possible contamination of outpost water supply.

Copyright © The Mars Society. All rights reserved. | Main Site